Diosmetin is a natural flavonoid present in the rind of ‘kalamansi’ and the leaves of ‘alagao’. It is a good antioxidant and anti-inflammatory agent. It is medically prescribed to treat hemorrhoids and varicose veins. Its molecular weight is 300.26 g/mol.

(a) How many millimoles (mmol) of diosmetin are in a caplet that contains 250 mg of diosmetin?

(b) If one is prescribed to take 3 caplets a day, how many milligrams of diosmetin are to be taken a day?

(c) If one has to obtain 1.5 mole of diosmetin for treatment, how many caplets does he need?


1
Expert's answer
2020-10-12T13:31:02-0400

Molar mass of diosmetin "= 300.26\ g/mol"

"(a)" Millimoles in "250 \ mg" of diosmetin or one caplet"= \frac{250 \times 10^{-3}}{300.26}\times 10^3 =" "0.832"

"(b)" One caplet contain "250" "\ mg" of diosmetin.

Three tablets would contain "250\times 3= 750\ mg" of diosmetin.

"(c)" Required amount "=" 1  mole "= 1000" millimoles.

Millimoles in "1 \"caplet"= 0.832"

Number of caplets required for this amount"=\frac{1000}{0.832}=" "1201.92\approx 1202"